having trouble with the following
k-10/2 = -7
the answer given on worksheet is -4, but I get 3
can you explain for me.
thanks.

Answer by Alan3354(69443)

k-10/2 = -7
the answer given on worksheet is -4, but I get 3
---------------
(k-10)/2 = -7
Multiply by 2
k-10 = -14
Add 10
k = -4